Mart is 4 times the age of his son Matthew. In 8 years, he will be eight-thirds of his son's age. How old is Mart
1
Expert's answer
2020-05-25T13:49:37-0400

Mart = 4 * Matthew

Mart + 8 = (Matthew + 8) * 8 / 3

4* Matthew + 8 = 8 / 3 * Matthew + 64 / 3

4/3 * Matthew = 40 /3

Matthew = 10

Mart = 4 * 10 = 40